# Comparative Analysis of Weather-Based Indexes and the Actuaries Climate IndexTM for Crop Yield Prediction

### v1

@@ 4.1 Data
Annual crop yield data (in bushels per acre) for three major agricultural products-corn (grain), wheat (winter), and soybeans-produced in the U.S. is obtained from the United States Department of Agriculture (https://quickstats.nass.usda.gov/, last accessed on April 18, 2025) for all states, covering the period from 1961 to 2023. Following the same approach used in regional weather-based index calculations, sub-regional crop yields are calculated by averaging the annual yields of the states within each region. Analyses for corn and wheat are conducted across all regions, while for soybeans, due to data limitations, the analyses are restricted to four regions: CEA, CWP, MID, and SEA.
4.2 Results
@@
The analyses start with determining the dependence among three groups (Table 1). It is noticed that all correlation coefficients, $r$, are statistically significant ($p$-value $<0.01$), as shown in Table 3. Despite methodological differences in constructing the indexes, strong positive correlations are observed between the weather-based indexes and their corresponding components of the ACI. To assess the similarity of the distributions, we run K-S test which evaluates the null hypothesis that the distributions are identical. Table 3 also presents the indexes that exhibit statistically similar distributions, as indicated by $p$-values greater than 0.05. The results reveal that the precipitation-related indexes in Group-3 exhibit similar distributions across most regions compared to temperature measurements analyzed in the first two groups. Detailed descriptive statistics for each group are provided in Table B1 of the Appendix. Additionally, the indexes and their corresponding cumulative distribution function (CDF) values are illustrated in Figures B1 and B2 of the Appendix.
@@
Group-1 (CDD-T90) r 0.88 0.90 0.89 0.94 0.86 0.96
K-S test (p) 0.089 <0.01 <0.01 <0.01 <0.01 0.408
Group-2 (HDD-T10) r 0.89 0.83 0.86 0.87 0.81 0.79
K-S test (p) 0.011 <0.01 <0.01 0.019 0.204 0.055
Group-3 (PRE-P) r 0.83 0.81 0.87 0.87 0.68 0.85
K-S test (p) 0.293 0.293 0.408 0.089 <0.01 0.089
Note. r: Spearman rank correlation coefficient.
In the second stage of the analysis, explanatory variables in each model are transformed into their principal components using PCA and FPCA. Figure 2 displays the total variation in the data sets, represented as regional averages, explained by the first four principal components derived from both methods for each model. The results demonstrate that the total variance explained by FPCA consistently surpasses that of PCA across all models, highlighting FPCA’s superior performance when dealing with time-dependent data sets like climate indexes.
For a detailed comparison of ACI components and their WBI counterparts, it is important to recall the structure of models outlined in Table 2. The first seven models are based on WBI components, while the subsequent seven models (Models 8–14) use their ACI counterparts. Comparing the total explained variance for paired models (e.g., 1 vs. 8, 2 vs. 9, 3 vs. 10, and so forth), it is clear that these values are nearly equivalent. For instance, the total variance explained by PCA for Model 4 is 83%, whereas it is 84% for Model 11. Similarly, FPCA explains 88% of the variance for Model 7 and 90% for Model 14. Since the choice of explanatory variables directly influence regression outcomes, ensuring high and comparable explained variance values across both index groups is crucial. Given that FPCA consistently explains a higher proportion of variance across all models, with values exceeding 85%, the principal components derived from FPCA are selected for the subsequent analyses, ensuring better model accuracy and reliability.
The distributions from the exponential family that best fit each crop yield for each region are identified, and the results are presented in Table 4. These distributions form the basis for estimating crop yields using GLMs and GAMs. Moreover, the incorporation of region-specific distributions ensures that the regression models account for spatial variability in crop yields, further improving the precision and relevance of the analyses.
Table 4: Best-fitting distributions for each crop
Region Crop
Corn (Grain) Wheat (Winter) Soybeans
CEA gamma gamma gamma
CWP normal normal -
MID normal gamma normal
SEA gamma gamma gamma
SPL normal normal gamma
SWP normal normal -
[…]
Method Model
1 2 3 4 5 6 7 8 9 10 11
GLM 0.288 0.296 0.170 0.286 0.334 0.353 0.384 0.217 0.231 0.292 0.248
GAM 0.384 0.388 0.309 0.396 0.475 0.414 0.444 0.389 0.446 0.416 0.381
XGB 0.738 0.747 0.691 0.753 0.757 0.766 0.769 0.690 0.732 0.734 0.708
LGBM 0.327 0.309 0.219 0.352 0.349 0.406 0.416 0.263 0.329 0.339 0.315
12 13 14 15 16 17 18 19 20 21 22
GLM 0.372 0.391 0.406 0.392 0.378 0.574 0.420 0.584 0.551 0.515 0.403
GAM 0.481 0.553 0.454 0.457 0.533 0.644 0.469 0.689 0.677 0.625 0.516
XGB 0.775 0.797 0.806 0.780 0.790 0.858 0.790 0.866 0.864 0.856 0.788
LGBM 0.437 0.456 0.427 0.401 0.429 0.571 0.414 0.616 0.581 0.523 0.406
Figure 3: MAPE values (Test) of each model as regional averages, Corn (Grain)
[…]
Method Model
1 2 3 4 5 6 7 8 9 10 11
GLM 0.262 0.296 0.118 0.301 0.282 0.345 0.333 0.215 0.260 0.239 0.242
GAM 0.439 0.515 0.360 0.527 0.372 0.400 0.433 0.398 0.436 0.397 0.359
XGB 0.725 0.756 0.659 0.752 0.730 0.737 0.776 0.725 0.738 0.696 0.723
LGBM 0.289 0.353 0.162 0.346 0.300 0.321 0.320 0.328 0.318 0.260 0.288
12 13 14 15 16 17 18 19 20 21 22
GLM 0.360 0.349 0.355 0.397 0.333 0.526 0.366 0.552 0.532 0.494 0.439
GAM 0.493 0.529 0.458 0.516 0.483 0.599 0.439 0.619 0.593 0.538 0.535
XGB 0.781 0.741 0.747 0.786 0.765 0.842 0.786 0.842 0.841 0.825 0.788
LGBM 0.367 0.371 0.327 0.417 0.362 0.524 0.408 0.552 0.538 0.507 0.440
Figure 4: MAPE values (Test) of each model as regional averages, Wheat (Winter)
[…]
Method Model
1 2 3 4 5 6 7 8 9 10 11
GLM 0.250 0.265 0.254 0.234 0.377 0.387 0.442 0.128 0.182 0.319 0.164
GAM 0.327 0.350 0.421 0.352 0.491 0.426 0.534 0.302 0.515 0.431 0.318
XGB 0.675 0.723 0.706 0.689 0.774 0.773 0.784 0.666 0.727 0.732 0.663
LGBM 0.222 0.252 0.258 0.286 0.357 0.426 0.425 0.146 0.314 0.326 0.235
12 13 14 15 16 17 18 19 20 21 22
GLM 0.341 0.378 0.402 0.362 0.389 0.602 0.374 0.597 0.573 0.567 0.381
GAM 0.467 0.571 0.479 0.441 0.615 0.721 0.422 0.641 0.695 0.606 0.528
XGB 0.740 0.786 0.761 0.770 0.768 0.862 0.752 0.861 0.847 0.838 0.756
LGBM 0.367 0.436 0.354 0.368 0.387 0.558 0.348 0.537 0.510 0.526 0.343
Figure 5: MAPE values (Test) of each model as regional averages, Soybeans
When analyzing the best-performing models for the three crops (Model 19 for corn and wheat, Model 17 for soybeans), it becomes clear that, in addition to temperature and precipitation variables, sea-level change plays a significant role in predicting yields. One possible explanation for this is that variations in sea-level can impact crop production more broadly, particularly through mechanisms such as increased soil salinity, flooding, and the loss of agricultural land. Furthermore, wind appears to have a less pronounced effect on soybean yield compared to corn and wheat. This could be attributed to the relatively short stature of soybean plants, making them more resistant to wind-induced damage. In contrast, taller crops like corn and wheat may be more vulnerable to lodging [^1] Lodging is the bending over of the stems of crops which makes them very difficult to harvest. It usually happens due to the wind-related stress. and other wind-related stresses, which can negatively impact their overall yield.
[…]
Another critical observation is that the final model (Model 22), which uses the composite ACI by aggregating all six components into a single index, consistently underperforms compared to Model 21, which incorporates each component separately, across all crops and methods. These findings highlight that the composite ACI index alone fails to fully capture and explain the complex weather conditions affecting crop yields. This suggests that treating the individual components separately provides a more nuanced and effective modeling approach, ensuring greater accuracy in predicting yield variations.
The second comparison analysis is designed to evaluate the performance sensitivity of machine learning algorithms with respect to different input variables. Table 8 presents only the results of XGB for the second comparison analysis, as it outperformed LGBM in earlier analyses. In Table 8, XGB refers to the results obtained using principal components derived from climate indexes, while XGB-C corresponds to the scenario where the climate indexes are directly used as input variables. Notably, with the XGB-C method, Model 21 (which incorporates all six components of ACI and thus constitutes the largest data set) consistently yields the best performance, with only a slight difference compared to the second-best models. Among the remaining models that exclude one or more components, the second-best performing models remain consistent with those from the first comparison analysis (Model 19 for corn and wheat, Model 17 for soybeans). One possible explanation for this is that the direct use of climate indexes preserves all original information in the data sets, enabling machine learning algorithms to process and use a greater number of features more effectively, compared to the more compressed data sets derived through principal component analysis. Regional performance metrics of XGB-C are given in Tables B11B11 of the Appendix.
It is also noteworthy that across all analyses, the best-performing models for each crop remain consistent across all methods (GLM, GAM, XGB, LGBM). This consistency underscores the substantial impact of specific climate indexes on crop yields, demonstrating that their predictive value holds true regardless of the modeling approach.
Table 8: Performance Comparison-2 as regional averages for each crop
Corn (Grain) Wheat (Winter) Soybeans
pR2 (Train) MAPE (Test) pR2 (Train) MAPE (Test) pR2 (Train) MAPE (Test)
Model XGB XGB-C XGB XGB-C XGB XGB-C XGB XGB-C XGB XGB-C XGB XGB-C
1 0.738 0.741 0.218 0.199 0.725 0.704 0.170 0.166 0.675 0.713 0.213 0.226
2 0.747 0.723 0.240 0.243 0.756 0.746 0.226 0.197 0.723 0.703 0.260 0.276
3 0.691 0.663 0.321 0.323 0.659 0.660 0.269 0.267 0.706 0.723 0.289 0.285
4 0.753 0.822 0.226 0.221 0.752 0.814 0.200 0.187 0.689 0.819 0.247 0.264
5 0.757 0.848 0.204 0.207 0.730 0.822 0.200 0.190 0.774 0.834 0.201 0.222
6 0.766 0.842 0.263 0.246 0.737 0.849 0.217 0.193 0.773 0.859 0.268 0.260
7 0.769 0.877 0.212 0.208 0.776 0.881 0.184 0.184 0.784 0.886 0.205 0.233
8 0.690 0.726 0.242 0.260 0.725 0.718 0.220 0.209 0.666 0.630 0.273 0.294
9 0.732 0.728 0.239 0.235 0.738 0.755 0.214 0.202 0.727 0.731 0.258 0.253
10 0.734 0.750 0.310 0.326 0.696 0.740 0.227 0.243 0.732 0.750 0.287 0.303
11 0.708 0.811 0.256 0.218 0.723 0.811 0.246 0.208 0.663 0.776 0.265 0.236
12 0.775 0.880 0.224 0.235 0.781 0.852 0.208 0.201 0.740 0.845 0.250 0.260
13 0.797 0.853 0.217 0.225 0.741 0.859 0.209 0.186 0.786 0.846 0.206 0.245
14 0.806 0.894 0.243 0.206 0.747 0.885 0.237 0.187 0.761 0.879 0.274 0.234
15 0.780 0.933 0.196 0.179 0.786 0.922 0.161 0.142 0.770 0.927 0.214 0.209
16 0.790 0.909 0.261 0.211 0.765 0.899 0.232 0.193 0.768 0.893 0.249 0.230
17 0.858 0.950 0.145 0.141 0.842 0.943 0.152 0.132 0.862 0.957ϕ 0.134 0.134ϕ
18 0.790 0.941 0.201 0.180 0.786 0.935 0.192 0.138 0.752 0.935 0.228 0.197
19 0.866 0.963ϕ 0.143 0.136ϕ 0.842 0.954ϕ 0.146 0.124ϕ 0.861 0.941 0.142 0.134
20 0.864 0.961 0.161 0.144 0.841 0.949 0.184 0.133 0.847 0.953 0.139 0.140
21 0.856 0.964 0.150 0.135 0.825 0.960 0.161 0.119 0.838 0.963 0.148 0.130
22 0.788 0.780 0.203 0.203 0.788 0.773 0.197 0.166 0.756 0.772 0.249 0.228
Note. Best results, ϕSecond best results.

### v2

@@ 4 Empirical Analysis
Annual crop yield data (in bushels per acre) for three major agricultural products-corn (grain), wheat (winter), and soybeans-produced in the U.S. is obtained from USDA’s National Agricultural Statistical Service (NASS) Quick Stats portal [^3] NASS Quick Stats: https://quickstats.nass.usda.gov/ for all states, covering the period from 1961 to 2024. Following the same approach used in regional weather-based index calculations, sub-regional crop yields are calculated by averaging the annual yields of the states within each region. Analyses for corn and wheat are conducted across all regions, while for soybeans, due to data limitations, the analyses are restricted to four regions: CEA, CWP, MID, and SEA.
The climate impacts of weather-based indexes and ACI components on the agricultural and financial markets are analyzed using data from six U.S. regions, and the corresponding findings are presented.
4.1 Agricultural Case
Firstly, the explanatory variables in each model (Table 1) are transformed into their principal components using PCA and FPCA. Figure 3 illustrates the total variation in the data sets, represented as regional averages, explained by the first four principal components derived from both methods for each model. The results demonstrate that the total variance explained by FPCA consistently surpasses that of PCA across all models, highlighting FPCA’s superior performance when dealing with time-dependent data sets like climate indexes.
For a detailed comparison, recall that the first seven models are constructed based on WBIs, while the subsequent seven models (Models 8-14) use their ACI counterparts. When comparing the total explained variance between these paired models (e.g., 1 vs. 8, 2 vs. 9, 3 vs. 10, and so forth), it is clear that the values are nearly equivalent. For instance, PCA explains 70% of the variance in Model 5, whereas it explains 71% in Model 12. Similarly, FPCA captures 88% of the variance in Model 7 and 90% in Model 14. Since the choice of explanatory variables, and consequently the derived principal components, directly affects the outcomes of the analyses, ensuring high and comparable explained variance values across both index groups is crucial for robustness. Given that FPCA consistently accounts for a larger proportion of variance across all models, with values exceeding 85%, the FPCA-derived principal components are selected for the subsequent analyses, ensuring better model accuracy and reliability.
Secondly, potential time trends in the crop yields are examined using the ADF test, and where necessary, the yields are detrended using Equation (9). The regional ADF test results for all crops are summarized in Table 3, where $p$-values below 0.05 indicate that the data are time-stationary. As shown in Table 3, only the SPL and SWP regions for corn are initially found to be stationary; therefore, no transformation is applied to the data in these regions. The repeated ADF test results obtained after detrending confirm that all crop yields become time-stationary. The detrended crop yields, illustrated for each region in Figure 4, are then used as the dependent variables in the regression and machine learning analyses.
Table 3: Regional ADF test results for each crop
ADF test
($p$-value)
Crop Region
CEA CWP MID SEA SPL SWP
Before
Detrend
Corn 0.997 0.787 0.971 0.902 0.045 0.033
Wheat 0.959 0.314 0.996 0.996 0.546 0.181
Soybeans 0.717 - 0.993 0.961 0.994 -
After
Detrend
Corn <0.01 0.049 <0.01 <0.01 - -
Wheat <0.01 <0.01 0.038 <0.01 <0.01 0.047
Soybeans <0.01 - <0.01 0.041 <0.01 -
Figure 4: Detrended crop yields over time
Thirdly, to establish a solid foundation for crop yields using GLMs and GAMs, the best-fitting distribution from the exponential family is identified using detrended crop yields for all regions. The resulting distributions are presented in Table 4. The incorporation of region-specific distributions ensures that the regression models account for spatial variability in crop yields, further improving the precision and relevance of the analyses.
Table 4: Best-fitting distributions for each crop
Region Crop
Corn Wheat Soybeans
CEA gamma normal gamma
CWP gamma normal -
MID gamma gamma gamma
SEA normal normal gamma
SPL normal gamma gamma
SWP normal normal -
Before obtaining the performance results of the models, the parameters for cross-validation are set as $N=64$, $M=5$, and $k=6$ in Equations (13) and (14). The MAPE values as regional averages for the training data of each crop are presented in Table 5. Based on the average values presented in the last row of Table 5, the GAM performs slightly better than the GLM, achieving lower MAPE values, though the difference remains relatively small. A similar pattern is observed between the XGB and LGBM, where XGB consistently outperforms LGBM, ultimately producing the lowest MAPE values among all methods. For the training data set, the average error rates for XGB range between 4% and 6% across all crops, whereas for LGBM, this range increases to 8%-15%. In contrast, the regression-based models exhibit significantly higher error rates, approximately around 30%.
Table 5: MAPE values as regional averages for each crop, model and method (Train set)
Model Corn Wheat Soybeans
GLM GAM XGB LGBM GLM GAM XGB LGBM GLM GAM XGB LGBM
1 0.362 0.345 0.071 0.154 0.282 0.278 0.047 0.089 0.297 0.298 0.039 0.084
2 0.364 0.358 0.080 0.161 0.283 0.276 0.047 0.089 0.295 0.296 0.049 0.090
3 0.383 0.374 0.079 0.166 0.283 0.281 0.042 0.086 0.295 0.296 0.038 0.084
4 0.359 0.357 0.075 0.153 0.282 0.279 0.048 0.088 0.295 0.296 0.043 0.086
5 0.365 0.345 0.071 0.157 0.282 0.278 0.044 0.086 0.294 0.295 0.035 0.082
6 0.356 0.340 0.064 0.154 0.282 0.281 0.040 0.083 0.294 0.293 0.038 0.086
7 0.349 0.348 0.069 0.152 0.281 0.277 0.040 0.084 0.293 0.293 0.039 0.084
8 0.368 0.351 0.075 0.153 0.285 0.283 0.047 0.089 0.296 0.297 0.041 0.085
9 0.368 0.355 0.077 0.161 0.283 0.281 0.048 0.089 0.297 0.296 0.045 0.087
10 0.369 0.352 0.074 0.161 0.280 0.280 0.043 0.085 0.294 0.292 0.043 0.089
11 0.364 0.336 0.072 0.155 0.283 0.282 0.048 0.088 0.295 0.295 0.043 0.086
12 0.357 0.345 0.066 0.148 0.278 0.275 0.039 0.084 0.294 0.294 0.041 0.084
13 0.353 0.347 0.066 0.147 0.279 0.277 0.040 0.083 0.294 0.295 0.046 0.086
14 0.350 0.339 0.063 0.150 0.279 0.275 0.040 0.085 0.294 0.294 0.044 0.088
15 0.356 0.350 0.070 0.150 0.279 0.277 0.043 0.085 0.294 0.293 0.042 0.088
16 0.347 0.341 0.062 0.149 0.280 0.276 0.038 0.083 0.292 0.292 0.045 0.089
17 0.336 0.325 0.054 0.130 0.280 0.280 0.040 0.084 0.293 0.293 0.043 0.086
18 0.351 0.340 0.068 0.153 0.280 0.277 0.041 0.082 0.292 0.292 0.045 0.088
19 0.335 0.312 0.052 0.138 0.279 0.280 0.043 0.085 0.294 0.294 0.043 0.087
20 0.332 0.314 0.050 0.137 0.280 0.278 0.040 0.084 0.292 0.291 0.040 0.086
21 0.336 0.314 0.054 0.143 0.279 0.276 0.044 0.085 0.291 0.290 0.044 0.087
22 0.357 0.342 0.071 0.151 0.285 0.284 0.045 0.089 0.293 0.292 0.049 0.091
Avg. 0.355 0.342 0.067 0.151 0.281 0.279 0.043 0.086 0.294 0.294 0.042 0.087
Since the GLM and GAM yield unsatisfactory results for the training set and XGM consistently outperforms LGBM, the MAPE values as regional averages for the test set are reported only for XGB in Figure 5. However, for a more comprehensive comparison across models, the regional MAPE values computed separately for both training and test data sets for all crops are given in Tables B2-B5 of the Appendix.
For corn, Model 19 achieves the best performance with a MAPE of 8.3%. The results of each model facilitate a detailed examination of how individual ACI components influence crop yields. For instance, Model 14, which includes temperature and precipitation components (T90, T10, P), achieves a MAPE of 10.4%. Sequentially adding the wind (W), drought (D), and sea-level (S) components in Models 15, 16, and 17, respectively, allows to evaluate their incremental effects. Among these, Model 17 produces the best result with a MAPE of 8.4%, indicating that the sea-level component has the strongest positive effect on corn yield prediction, reducing the error from 10.4% to 8.4%. Furthermore, when the wind and sea-level components are combined (Model 19), model performance improves slightly further, reaching 8.3% MAPE. Interestingly, Model 21, which incorporates all six ACI components, performs worse than Model 19. This outcome suggests that while wind and sea-level components enhance predictive accuracy, the inclusion of the drought has a diminishing effect on model performance when all other five climate variables of ACI are already used as predictors.
For wheat, similar to the findings for corn, Model 19 achieves the best overall performance, yielding a MAPE of 8.2%. When examining the individual component models, Model 8 (T90), Model 9 (T10), and Model 10 (P) produce MAPE values of 9.2%, 8.7%, and 8.4%, respectively. These results indicate that, when analyzed individually, precipitation has a relatively higher impact on wheat yield prediction than the temperature-related components. When precipitation is combined with low-temperature extremes (as in Model 13), model’s performance improves slightly, with the MAPE decreasing to 8.3%. Furthermore, considering Models 15 (9.2%), 16 (8.8%), and 17 (8.7%), which sequentially incorporate wind, drought, and sea-level components, alongside temperature and precipitation, the results reveal that the sea-level component contributes the most substantial improvement in predictive accuracy.
For soybeans, the results differ slightly from those obtained for corn and wheat. The best-performing model for soybeans is Model 17 (MAPE of 7.5%). Unlike the findings for the other two crops, the inclusion of the wind component appears to negatively affect model performance for soybean yield. The performance difference between Model 12 (T90 and P) and Model 13 (T10 and P) is noteworthy. Model 12 achieves a MAPE of 8.2%, while Model 13 yields a higher error of 8.9%, indicating that warm weather extremes (T90) appear to have a stronger influence on soybean yield than cold weather extremes (T10). This conclusion is further supported by the performance results of Model 8 (T90: 8.4%) and Model 9 (T10: 9.2%).
Figure 5: MAPE values as regional averages for each crop and model with XGB (Test set)
When analyzing the best-performing models for the three crops (Model 19 for corn and wheat, Model 17 for soybeans), it becomes clear that, alongside the temperature and precipitation components, sea-level change plays a significant role in yield prediction. One possible explanation for this is that variations in sea-level can impact crop productivity more broadly through mechanisms such as increased soil salinity, flooding, and the loss of agricultural land. Furthermore, wind appears to have a less pronounced effect on soybean yield compared to corn and wheat, likely due to the relatively short stature of soybean plants, which makes them more resistant to wind-induced damage. In contrast, taller crops like corn and wheat are more vulnerable to lodging [^4] Lodging refers to the bending or collapse of crop stems, often caused by strong winds, which makes harvesting difficult. and other wind-related stresses that can negatively impact their overall yield.
A detailed comparison of the performance results for paired models (1 vs. 8, 2 vs. 9, 3 vs. 10, and so forth) on crop yields also reveals consistent and satisfactory outcomes. For instance, the performance difference between Model 1 (CDD: 11.3%) and Model 8 (T90: 11.4%) for corn, Model 2 (HDD: 9.2%) and Model 9 (T10: 8.7%) for wheat, Model 3 (PRE: 7.7%) and Model 10 (P: 8.2%) for soybeans are almost negligible, suggesting comparable explanatory power between ACI components and their WBI counterparts.
Another critical observation is that Model 22, which uses the composite ACI by aggregating all six components into a single index, consistently underperforms compared to Model 21, which incorporates each component separately, across all crops. These results emphasize that the composite ACI alone fails to fully capture and explain the complex weather conditions affecting crop yields. Consequently, analyzing the individual components separately provides a more nuanced and effective modeling approach, leading to improved accuracy in predicting yield variations.
The second comparison in the analysis (see Figure 1) is designed to evaluate the performance sensitivity of machine learning algorithms with respect to different input variables. In Table 6, XGB refers to the results obtained using principal components derived from climate indexes, while XGB-C corresponds to the scenario where the climate indexes are directly used as input variables. Notably, with the XGB-C method, Model 21, which constitutes the largest data set, consistently yields the best performance, with only a slight difference compared to the second-best models. Among the remaining models that exclude one or more components, the second-best performing models remain consistent with those from the first comparison analysis (Model 19 for corn and wheat, Model 17 for soybeans). One possible explanation for this is that the direct use of climate indexes preserves all original information in the data sets, enabling machine learning algorithms to process and use a greater number of features more effectively, compared to the more compressed data sets derived through principal component analysis. Regional performance metrics of XGB-C are also given in Table B6 of the Appendix.
Table 6: Performance Comparison-2: MAPE values as regional averages for each crop (Test set)
Corn Wheat Soybeans
Model XGB XGB-C XGB XGB-C XGB XGB-C
1 0.113 0.103 0.093 0.092 0.086 0.085
2 0.127 0.134 0.092 0.091 0.100 0.100
3 0.143 0.149 0.087 0.086 0.077 0.072
4 0.107 0.107 0.095 0.090 0.093 0.086
5 0.114 0.103 0.093 0.086 0.077 0.075
6 0.118 0.125 0.087 0.084 0.082 0.073
7 0.111 0.106 0.087 0.085 0.083 0.077
8 0.114 0.102 0.092 0.089 0.084 0.087
9 0.124 0.126 0.087 0.088 0.092 0.099
10 0.140 0.140 0.084 0.085 0.082 0.082
11 0.106 0.106 0.087 0.085 0.088 0.089
12 0.110 0.105 0.091 0.084 0.082 0.083
13 0.113 0.113 0.083 0.087 0.089 0.085
14 0.104 0.105 0.089 0.084 0.081 0.082
15 0.118 0.110 0.092 0.087 0.081 0.080
16 0.120 0.115 0.088 0.086 0.083 0.084
17 0.084 0.087 0.087 0.085 0.075 0.068ϕ
18 0.124 0.118 0.093 0.084 0.080 0.080
19 0.083 0.078ϕ 0.082 0.084ϕ 0.088 0.074
20 0.090 0.085 0.096 0.088 0.080 0.078
21 0.098 0.075 0.095 0.082 0.085 0.068
22 0.110 0.096 0.098 0.094 0.096 0.094
Note. Best results, ϕSecond best results.
4.2 Financial Market Case
@@
The analysis of weather derivatives pricing begins by examining the dependence among the three groups (see Table 2). Table 7 presents the correlation coefficients ($\rho$) between the weather-based indexes and their corresponding ACI components within each group, along with the $p$-values from the A-D test, which evaluates the null hypothesis that the two distributions are identical. Despite methodological differences in constructing the indexes, strong positive correlations are observed between the corresponding indexes. For the A-D test, $p$-values greater than 0.05 (bold in Table 7) indicate that the paired indexes exhibit statistically similar distributions. The results show that the precipitation-related indexes in Group-3 tend to display similar distributions across most regions compared to temperature-based indexes analyzed in the first two groups. Detailed descriptive statistics for each group are provided in Table B1 of the Appendix.
@@
Group-1
(CDD-T90)
$\rho$ 0.88 0.90 0.89 0.94 0.86 0.97
A-D test ($p$-value) 0.040 <0.01 <0.01 <0.01 <0.01 0.250
Group-2
(HDD-T10)
$\rho$ 0.89 0.83 0.86 0.87 0.81 0.80
A-D test ($p$-value) 0.025 <0.01 <0.01 <0.01 0.105 <0.01
Group-3
(PRE-P)
$\rho$ 0.83 0.79 0.88 0.87 0.69 0.84
A-D test ($p$-value) 0.250 0.126 0.250 0.189 <0.01 0.085
Note. $\rho$: Spearman rank correlation coefficient.
For the payoff analysis of derivative options, pairs of indexes with the highest and lowest correlations are selected as case studies. As shown in Table 7, the pair with the strongest correlation is identified in Group-1 for the SWP region ($\rho=0.97$), while the weakest correlation is found in Group-3 for the SPL region ($\rho=0.69$). Call option payoffs are calculated using Equation (15), based on the index values from these two groups and regions.
Additionally, the key assumption of the HBA and IM pricing methods, which is that the underlying index should be time-stationary, is also considered. According to the ADF test results (not shown), the CDD and T90 indexes in Group-1 for the SWP region are determined to be non-stationary; therefore, these series are detrended using Equation (9) prior to inclusion in the analyses.
For consistency, the tick size ($\alpha$) in the Equation (15) is set to 100 for both groups, and the strike level ($K$) is defined as the mean of the historical index values plus 20% of their standard deviation (Jewson and Brix 2005). In the HBA method, payoffs for each year are computed, and the correlations between the payoffs and their A-D test results are analyzed. For the IM method, the best-fitting distributions for the historical indexes are identified, and 1,000 index values are simulated from these distributions to compute payoffs. Finally, the A-D test is applied to the simulated payoffs values. The results of the similarity tests comparing the payoffs obtained from both methods are summarized in Table 8.
Table 8: Similarity tests between payoffs for each method
Method N Compared Parameters Group-1 (SWP) Group-3 (SPL)
CDD T90 PRE P
HBA 64 Payoffs $\rho$ 0.82 0.65
A-D test ($p$-value) 0.250 0.250
IM 64 Indexes Best-fitting dist. weibull beta lognorm weibull
1,000 Payoffs A-D test ($p$-value) <0.01 0.038
Note. $\rho$: Spearman rank correlation coefficient, N: Number of years.
According to Table 8, the correlations among the payoffs obtained using the HBA method closely align with the correlations observed between the original indexes. For Group-1 (SWP), the correlation coefficient between the payoffs is 0.82 (down from 0.97), while for Group-3 (SPL), it is 0.65 (decreasing from 0.69). The decline observed for Group-1 (SWP) is expected, as the CDD and T90 indexes are detrended prior to the analysis. Moreover, based on the A-D test results for the HBA method ($p>0.05$), the payoff distributions are statistically similar. Interestingly, although the A-D test results for Group-3 (SPL) in Table 7 show no distributional similarity between the indexes themselves, the observed similarity among their payoffs implies that these indexes exhibit consistent behavior in terms of financial outcomes, highlighting their potential applicability in weather-derivative markets.
Table 8 reports the best-fitting distributions for each index pair (with detrended CDD and T90). Distinct best-fitting distributions are identified for all indexes, a finding further confirmed by the A-D test results derived from 1,000 simulated indexes and their corresponding payoffs ($p<0.05$). In contrast to the HBA findings, no similar distributions are detected among the simulated payoffs in the IM approach, indicating that the degree of similarity among payoffs in simulation-based analyses strongly depends on the distributional characteristics of the underlying indexes themselves.
The payoffs and ECDF values derived from both methods are illustrated in Figure 6. Examination of the payoff graphs reveals that the call options constructed from the indexes within each group often generate payments during similar time periods. This pattern indicates that the indexes tend to reach extreme values simultaneously, confirming their consistency and suitability for use in weather-derivative markets.
The fair price of uncapped call option can be determined through the HBA method by averaging all historical payoffs. For Group-1, the fair price is calculated as 2,192.23 for CDD and 34.74 for T90. Similarly, for Group-3, the corresponding fair prices are 36.34 for PRE and 40.61 for P. The observed discrepancies in fair prices arise from the fact that each index expresses weather conditions using different measurement units. Recall that CDD quantifies temperature in degree-days, whereas PRE measures precipitation in millimeters. In contrast, the ACI components T90 and P express weather conditions as standardized deviations relative to a reference period. Comparing the prices obtained using T90 and P, which share identical units, indicates that the call option priced with P is more expensive than that based on T90 (40.61 vs. 34.74). This finding suggests that extreme precipitation events have historically occurred more frequently than extreme temperature events.
(a) Group-1 (SWP)
(b) Group-3 (SPL)
Figure 6: Payoffs and ECDF values for each method